OpenStudy (anonymous):

Find the particular solution for the anti-derivative of f'(x) = 2x^3 + 1 – sin x, if f(0) = 1.

OpenStudy (anonymous):

Take integral

OpenStudy (anonymous):

That doesn't help my confusion... :(

OpenStudy (anonymous):

Do you know how to take integral

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Can you teach me?

OpenStudy (anonymous):

Integration is a broad subject but here is the basic integration technique \[\int{ax^b}=\frac{a}{b+1}x^{b+1}\]

OpenStudy (anonymous):

Left out dx above \[\int{ax^b dx}=\frac{a}{b+1}x^{b+1}\]

OpenStudy (anonymous):

okay. still confuseddd